# Detecting Neurocognitive Disorders through Analyses of Topic Evolution and Cross-Modal Consistency in Visual-Stimulated Narratives

### v2

@@ III-B CU-MARVEL-RABBIT
The CU-MARVEL Corpus is designed and collected by our group in Hong Kong, encompassing various cognitive tests. To assess the cognitive capacities of individuals with NCDs in integrating sequential visual information and constructing coherent narratives, we adopt the Rabbit Story task of CU-MARVEL, denoted as “CU-MARVEL-RABBIT” Corpus.
@@
The corpus records Cantonese-speaking participants narrating a story about a girl finding a rabbit through 15 cartoon images, as illustrated in Fig. 1(b). These images generally depict 5 sequential scenes (or topics): (1) the discovery and subsequent loss of the rabbit (images 1-4), (2) an encounter with a squirrel and some ants (images 5-7), (3) a meeting with a goose (images 7-9), (4) a run-in with a cow (9-11), and (5) a reunion with the rabbit followed by a farewell wave to its family (images 12-15).
@@
These images are presented in a booklet, with each image displayed on a separate page. Participants will first review all 15 images to familiarize themselves with the content. Then, they will examine each image individually, providing a description of its content before proceeding to the next image. This process will continue sequentially until the stories for all 15 images have been completed.
Figure 2: Illustrative Cantonese narratives in CU-MARVEL-RABBIT, where English words are roughly translated by GPT-4o for reader’s comprehension only.
@@
The CU-MARVEL-RABBIT Corpus (version “2024-06-13”) consists of 758 subjects, including 467 HCs and 291 individuals with NCDs. Illustrative examples in Fig. 2, where the English translations are coarsely rendered by GPT-4o [71] for reader’s comprehension only, highlight the typical differences between HC and NCD narratives. Specifically, HC tends to capture most visual details (e.g., characters (girl, dog, rabbit, squirrel, goose, cow), objects (tree hollow, rock), and settings (house, garden, forest)) and weaves them into a clear, logical storyline across the images (e.g., rabbit was missing → girl and dog searched in the forest → rabbit was found with its family). In contrast, NCD often misses or misidentifies key details (e.g., squirrel, goose) and exhibits difficulties in establishing coherent connections between events (primarily isolated descriptions).
@@
The data distribution and train-test split are listed in Table I. To benchmark the performance of frontend tools (including speech recognition, word segmentation, and part-of-speech tagging), we extensively annotated a subset of 46 participants (22 HCs, 24 NCDs) from the test set (aligned with [48]), naming “Frontend-Evaluation” set. The CU-MARVEL-RABBIT Corpus records the assessment for each subject by specialists in neurology, using a 0–4 severity scale for cognitive impairment, with lower values corresponding to better cognitive performance. Specifically, values 0-1 indicate HCs and 2, 3, and 4 represent progressively severe NCDs.
@@ IV Approach
In this work, we propose a comprehensive analysis for the automatic detection of Neurocognitive Disorders based on VSNs, encompassing both microstructural and macrostructural analysis. Microstructural analysis, detailed in Section IV-A, focuses on the extraction of acoustic and linguistic features selected from previous literature. On the other hand, macrostructural analysis, presented in Sections IV-B to IV-D, introduces three distinct methodologies to capture global and temporal patterns with varying focal points.
@@ IV-A Traditional Acoustic and Linguistic Analysis
Based on the literature review, the microstructural features have been categorized into two main groups: acoustic features and linguistic features. We select 10 acoustic features and 13 linguistic features from the literature that are basic, universal and suitable for Cantonese in the CU-MARVEL-RABBIT Corpus, as summarized in Table II. Before extracting these features, we preprocess all audios with FullSubNet+ [72] denoiser and peak-normalization to ensure internal consistency and avoid the Clever Hans Effect (where the model exploits irrelevant or unintended shortcuts like noise) [73].
The rows of A1-A10 in Table II list the acoustic features, related to pause (A1-A6), speech duration (A7-A8), articulation rate (A9) and speaking rate (A10). The articulation rate measures phonetic production speed as the number of syllables per voiced duration (excluding pauses), while the speaking rate quantifies overall delivery speed as the number of syllables per total duration (including pauses), thereby reflecting real-time communication fluency [74].
To extract these features, we first employ Voice Activity Detection (VAD) using the pretrained Feedforward Sequential Memory Network, as described in the paper [75], to obtain the timestamps of the voiced segments. With these segment timestamps, a pause is detected by the silence gap between two adjacent voiced segments when the duration exceeds the duration of a syllable for this speaker, which is calculated by dividing the total voiced duration by the total number of syllables for each Cantonese participant.
The linguistic features are extracted from the automatic transcriptions of participants obtained using the pretrained Automatic Speech Recognition (ASR) model SenseVoice-Small [76], which demonstrates a Word Error Rate (WER) of 31.37% on the ADReSS test set and a Character Error Rate (CER) of 34.22% on the Frontend-Evaluation set of the CU-MARVEL-RABBIT corpus. The rows of L1-L13 in Table II show the name and the corresponding description of each feature. To extract these linguistic features from Cantonese transcripts, we utilized pretrained BERT base models for word segmentation [^3] https://huggingface.co/ckiplab/bert-base-chinese-ws and POS auto-tagging [^4] https://huggingface.co/ckiplab/bert-base-chinese-pos, with F1 scores of 82.69% and 85.11% respectively on the Frontend-Evaluation set of the CU-MARVEL-RABBIT corpus.
@@
Figure 3: The topic distribution for the HC group (a) and the NCD group (b), and the topic evolution for the HC group (c) and the NCD group (d), as obtained from DTM. (a-b) show the weights of each topic word (top-10) within the topic (from A to E) for the group. (c-d) display average topic proportions with curves and deviations with filled areas. The order of topics A to E is based on the positions of their first peaks in topic evolutions.
@@ IV-B Reference-based Narrative Alignment
A critical part of evaluating the narrative task, particularly within the Rabbit Story context, involves assessing the comprehensiveness of a subject’s narrative. To this end, we identify keywords associated with the task and measure the coverage in the subject’s narratives. Specifically, we rank words from the HCs group (excluding punctuation and stopwords) based on their relevance to visual stimuli. This relevance is quantified through cosine similarity between global image embeddings (from the global special token [CLS]) and text embeddings (via mean pooling), both extracted from the final projection layers of the visual and textual encoders in the Chinese-CLIP model [77], which is pretrained via a contrastive learning paradigm that aligns images with corresponding textual descriptions on diverse Chinese variations (including the Cantonese dialect). As the textual and visual modalities are projected into a shared multimodal embedding space, the Chinese-CLIP model inherently enables cross-modal similarity measurement. These derived top-ranked words are categorized into four semantic groups informed by human prior knowledge of the Rabbit Story, and the Narrative Scoring Scheme (NSS) proposed by [24], including general settings or introduction (e.g., daytime/one day, garden/house/forest), characters or roles (e.g., girl, dog, rabbit), objects (e.g., sofa, box, stone), and actions (e.g., play, find, walk). Coverage (abbreviated as “cvg”, ranging from 0 to 1) for each category is then computed as features for analysis, focusing on the narrative content instead of linguistic form like PoS ratios. For example, the Rabbit Story features 7 characters, but the NCD narrative in Fig. 2(b) only mentions 4 of them, resulting in a “cvg_role” value of 4/7.
@@
While these coverage metrics effectively capture whether local narrative elements are addressed, they lack insights into global coherence and connections across the entire picture set. To address this limitation, metrics commonly used in image captioning tasks, such as BLEU-n (n=1,2,3,4), METEOR, and ROUGE-L, are employed to quantify the degree of alignment between subject narratives and reference descriptions. For this purpose, we engaged six research team members, all thoroughly familiar with the NSS rubric [^5] The NSS rubric evaluates narratives across seven dimensions: introduction, character development, mental and emotional states, referencing, conflict/resolution (or event/reaction), cohesion, and conclusion. Each dimension uses a 0-5 scale, with concrete elaborations and examples provided for the “Proficient (5)”, “Emerging (3)”, and “Minimal/Immature (1)” levels., to craft reference descriptions of the Rabbit Story that exemplify the proficient criteria. The final BLEU-n, METEOR, and ROUGE-L scores are averaged across these references for each subject, providing additional features that reflect global narrative coherence.
@@
In short, we derive two types of features for reference-based analysis: (1) 10 domain-specific coverage metrics, including introduction, character, object, and action words, as well as high-frequency visual-related terms (nouns, verbs, pronouns, adjectives, adverbs, and all types of POS); and (2) 6 expert-aligned narrative similarity scores (BLEU-n, METEOR, ROUGE-L). These reference-based features enable macrostructural analysis by evaluating informativeness and coherence against curated standards. However, as these features derive primarily from word frequencies or local n-gram matching, they are incapable of modeling temporal topic development or the consistency.
@@ IV-C DTM-based Temporal Analysis
Temporal changes or thematic evolution in narratives are critical for understanding the progression and manifestations of NCDs, as these elements reflect higher-level cognitive abilities. To capture these temporal dynamics, we utilize Dynamic Topic Modeling (DTM) [78], which tracks them by explicitly modeling the evolution of topics over time. In contrast to previously mentioned static methods, DTM offers a dynamic perspective on the underlying temporal trends within the narratives.
@@
Dynamic Topic Model (DTM) is an extension of classical topic models. It introduces time dependencies upon the distributions, making it particularly suitable for analyzing temporal changes in narrative content. Formally, DTM models the dynamic nature of topic evolution by introducing time dependency $t$ in topic word distribution $\beta_{k,t}$, where $k$ and $t$ denote the topic and time indices respectively. In this way, it allows the topic distributions to drift over time via a Gaussian process, i.e., $\beta_{t}|\beta_{t-1}\sim\mathcal{N}(\beta_{t-1},\boldsymbol{\Sigma})$, where $\boldsymbol{\Sigma}$ represents the covariance matrix controlling topic transition smoothness.
Then the generative process of DTM proceeds as follows:
Firstly, for each document $d$ in time slice $t$, the model first samples a topic proportion from a Dirichlet prior: $\theta_{d}\sim\text{Dir}(\alpha)$, where $\alpha$ is the Dirichlet concentration parameter controlling the sparsity of topic distributions.
Secondly, for each word position $n$ in document $d$, the model samples a topic assignment from a multinomial distribution parameterized by the document topic proportion: $z_{d,n}\sim\text{Mult}(\theta_{d})$.
@@
The model parameters could be estimated through variational inference [79] or the Expectation-Maximization [80] algorithms during the training phase.
@@
In order to apply the DTM model to narrative analysis for NCD detection, we evenly split each participant’s ASR transcription into 15 slices, assuming that each image leads to approximately similar number of words in spoken narratives. To mitigate fragmented sentences, we shift the split points to the nearest punctuation boundaries. This segmentation allows us to track the evolution of participants’ narrative topics across the story’s progression, enabling a granular analysis of discourse cohesion by observing how thematic content shifts and develops with each sequential image in the narrative.
@@
We use DTM by inputting the segmented transcription data, where each slice represents a distinct portion of the narrative associated with a time step. The DTM processes this data to identify evolving topics within and across the segmented slices, allowing us to observe how smoothly and coherently topics transition as the story progresses.
Figure 4: (a) The proposed model TITAN. (b) The Attention module of TITAN. (c) The correlation map between image and text embeddings in TITAN, where the 15 textual chunks are manually segmented and aligned to the corresponding images. TITAN takes images and text narratives as inputs to predict NCDs, with Attention using RoPE to focus on important positions.
@@
Fig. 3 illustrates the topic distribution (a, b) and temporal evolution of topic proportions (c, d) for the HC and NCD groups. For the distribution of topics over words, there exists differences between HC and NCD groups, where HC topics over time tend to be more aligned with the order of the images. For example, the topic words of the HC group, shown in Fig. 3(a), include “home” and “garden” in topic “A” (corresponds to settings of images 1-4), “ant”, “tree hollow” and “goose” in topic “B” (corresponds to images 5-8), “cow” and “stone” in topic “C” (corresponds to images 9-10), “fall” and “pond” in topic “D” (corresponds to images 10-11), and “many” (rabbits) and “mother” in topic “E” (corresponds to images 12-15). These topic words generally align with the different scenes along the story. In contrast, the topic words of the NCD group, shown in Fig. 3(b), are more chaotic. For example, “tree hollow” in topic “B” and “ant/squirrel” in topic “C” both predominantly appear in the same scene in images 5 and 6, indicating overlapping topic assignments.
@@
Another interesting finding is that the HC narratives exhibit a cyclical pattern. Specifically, most HC individuals mention “return” and “home” in topic ”E”, even though these elements are not depicted in the final images. This implies that HCs can tie the ending of the story back to its beginning, completing the girl’s quest to find the lost rabbit and bringing the story to a satisfying closure.
@@
For the dynamic proportions of these topics over time, as shown in Fig. 3(c, d), it is observable that the HC group exhibits more distinct peaks (with each topic dominating consecutive slices) and smoother transitions between topics over time (peaks rising sharply after the preceding topic decreases), while the NCD group shows more erratic changes with less pronounced peaks (e.g., reduced peak amplitudes and sporadic dominance).
@@
Based on these findings, we proposed 6 DTM-related statistics for NCD detection. The first metric, termed “topic_consistency”, is derived from the hit rate of the top-10 topic words each time, which are derived from a DTM pretrained on all narratives. This statistic measures how consistently the subject’s topic words align with the top-10 words over time, thus providing insights into the stability and coherence of topics across different groups (HC vs. NCD). For example, if the words in text slice include all top-10 topic words at the same timestamp $t$, then topic_consistency($t$) = 1.0. Higher topic consistency suggests a more stable thematic structure, indicating a potentially healthier cognitive state. The second metric, termed “topic_cycle”, which is simply a boolean value indicating whether the word “home” appears in the latter half of each narrative.
@@
The other four metrics are derived from the dynamic properties of topic evolution, including standard deviation (“topic_variability”), correlation (“topic_temporal_corr”) of the proportions over time, peak-to-peak range (difference between the highest and lowest values, termed as “topic_ptp_range”), and rate of change (averaged Euclidean distance of adjacent topic distributions over time, termed as “topic_change_rate”). These statistics capture topic development of VSNs and could be used as features followed by a downstream classifier for NCD detection, which are interpretable through their well-defined mathematical properties.

### v3

@@ III-B The CU-MARVEL-RABBIT-RABBIT Corpus
The CU-MARVEL-RABBIT corpus was designed and collected by our research group in Hong Kong as part of the broader CU-MARVEL project, which includes multiple cognitive assessment tasks. To evaluate the ability of individuals with NCDs to integrate sequential visual information and generate coherent narratives, we utilize the Rabbit Story task within the CU-MARVEL framework, referred to herein as the “CU-MARVEL-RABBIT” corpus.
@@
This corpus contains narratives from Cantonese-speaking participants telling a story illustrated through 15 cartoon images about a girl finding a rabbit (similar to “Frog, Where Are You?” [71]). These images generally depict 5 sequential scenes (or topics): (1) the discovery and subsequent loss of the rabbit (images 1-4), (2) an encounter with a squirrel and some ants (images 5-7), (3) a meeting with a goose (images 7-9), (4) a run-in with a cow (9-11), and (5) a reunion with the rabbit followed by a farewell wave to its family (images 12-15). Note that overlapping images (e.g., images 7 and 9) indicate transitional events, such as the emergence of the goose and the cow.
@@
These images are presented in a booklet, with each image displayed on a separate page. To ensure uniformity and comparability, all participants first review all 15 images to familiarize themselves with the content. Then, they examine each image individually, describing its content before moving on to the next. This process continues sequentially until narratives for all 15 images have been completed.
@@
Illustrative examples are provided in the Appendix. The participant in the HC group (Fig. 6) tends to capture most visual details (e.g., characters (girl, dog, rabbit, squirrel, ants, goose, cow); objects (tree hollow, rock); and settings (house, garden, forest)) and weaves them into a clear, logical storyline across the images (e.g., rabbit was missing → girl and dog searched in the forest → rabbit was found with its family). In contrast, the participant with NCDs (Fig. 7) often misses or misidentifies key details (e.g., squirrel, goose) and exhibits difficulties in establishing coherent connections between events, resulting primarily in isolated descriptions.
[…]
@@
The CU-MARVEL-RABBIT Corpus also records the specialist neurological assessment for each participant, using a 0-4 severity scale for cognitive impairments, with lower values corresponding to better cognitive performance. Specifically, values 0-1 indicate HCs, while 2, 3, and 4 represent progressively severe NCDs. Detailed data distribution and train-test split are listed in Table I.
@@ IV Approach
In this work, we propose a comprehensive framework for the automatic detection of NCDs using VSNs, encompassing both microstructural and macrostructural analyses. Microstructural analysis, detailed in Section IV-A, focuses on the extraction of acoustic and linguistic features based on previous studies. In contrast, macrostructural analysis, presented in Sections IV-B through IV-D, introduces three distinct methodologies aimed at capturing global and temporal narrative patterns.
@@ IV-A Traditional Acoustic and Linguistic Analysis
Microstructural features can be broadly categorized into two main groups: acoustic features and linguistic features. Based on prior studies, we select 10 acoustic features and 13 linguistic features that are basic, universal and suitable for Cantonese in the CU-MARVEL-RABBIT Corpus, as summarized in Table II. Before extracting these features, we preprocess all speech recordings with FullSubNet+ [73] denoiser and peak-normalization to ensure internal consistency and mitigate the risk of models learning spurious cues (i.e., the Clever Hans Effect [74]).
Rows of A1-A10 in Table II list the acoustic features, related to pause (A1-A6), speech duration (A7-A8), articulation rate (A9) and speaking rate (A10). The articulation rate measures phonetic production speed as the number of syllables per voiced duration (excluding pauses), while the speaking rate quantifies overall delivery speed as the number of syllables per total duration (including pauses), thereby reflecting real-time communication fluency [75].
To extract these features, we first employ Voice Activity Detection (VAD) using the pretrained Feedforward Sequential Memory Network, as described in the paper [76], to obtain the timestamps of the voiced segments. With these segment timestamps, a pause is detected as the silence gap between two adjacent voiced segments when the duration exceeds the average syllable duration for that speaker, which is calculated by dividing the total voiced duration by the total number of syllables for each Cantonese participant.
The linguistic features are extracted from the automatic transcripts obtained using the pretrained Automatic Speech Recognition (ASR) model SenseVoice-Small [77], which achieved a Word Error Rate (WER) of 31.37% on the ADReSS test set and a Character Error Rate (CER) of 34.22% on the CMR-Frontend-Evaluation set of the CU-MARVEL-RABBIT corpus. Rows of L1-L13 in Table II show the name and corresponding description of each feature. To extract these linguistic features from the Cantonese transcripts, we used pretrained BERT-based models for word segmentation [^3] https://huggingface.co/ckiplab/bert-base-chinese-ws and PoS auto-tagging [^4] https://huggingface.co/ckiplab/bert-base-chinese-pos, achieving F1 scores of 82.69% and 85.11% for the CMR-Frontend-Evaluation set, respectively.
@@ IV-B Reference-based Narrative Alignment
A critical part of evaluating the narrative task, particularly within the Rabbit Story context, involves assessing the comprehensiveness of a participant’s narrative. To this end, we identify task-related keywords and measure their coverage in the narratives. Specifically, we rank words from the HCs group (excluding punctuation and stopwords) based on their relevance to visual stimuli. This relevance is quantified using cosine similarity between global image embeddings (from the [CLS] token) and text embeddings (via mean pooling), both extracted from the final projection layers of the visual and textual encoders in the CLIP-like models (Chinese-CLIP [^5] https://huggingface.co/OFA-Sys/chinese-clip-vit-huge-patch14 [78] for CU-MARVEL-RABBIT corpus, and Multilingual-CLIP [^6] https://huggingface.co/M-CLIP/XLM-Roberta-Large-Vit-B-16Plus [79] for ADReSS and ADReSS corpora), which are pretrained via a contrastive learning paradigm that aligns images with corresponding textual descriptions. As the textual and visual modalities are projected into a shared multimodal embedding space, the Chinese-CLIP model inherently enables cross-modal similarity measurement. These derived top-ranked words are categorized into four semantic groups informed by human prior knowledge of the Rabbit Story, and the Narrative Scoring Scheme (NSS) [24], including general settings or introduction (e.g., daytime, one day; garden, house, forest), characters or roles (e.g., girl, dog, rabbit), objects (e.g., sofa, box, stone), and actions (e.g., play, find, walk). Coverage (abbreviated as “cvg”, ranging from 0 to 1) for each category is then computed as a feature for analysis, focusing on the narrative content rather than linguistic form (e.g., PoS ratios).
@@
While these coverage metrics effectively capture whether local narrative elements are addressed, they provide limited insights into global coherence and connections across the entire picture set. To address this limitation, metrics commonly used in image captioning tasks, such as BLEU-n (n=1, 2, 3, 4), METEOR, and ROUGE-L, are adopted to quantify the degree of alignment between participants’ narratives and reference descriptions. For this purpose, we engaged six research team members, all thoroughly familiar with the NSS rubric [^7] The NSS rubric evaluates narratives across seven dimensions: introduction, character development, mental and emotional states, referencing, conflict/resolution (or event/reaction), cohesion, and conclusion. Each dimension is rated on a 0-5 scale, with detailed elaborations and examples provided for the “Proficient (5)”, “Emerging (3)”, and “Minimal/Immature (1)” levels., to craft reference descriptions of the Rabbit Story that exemplify the proficient criteria. The final BLEU-n, METEOR, and ROUGE-L scores are averaged across these references for each participant, providing additional features that reflect global narrative coherence.
@@
In short, we derive two types of features for reference-based analysis: (1) 10 domain-specific coverage metrics, including introduction, character, object, and action words, as well as high-frequency visual-related terms (including nouns, verbs, pronouns, adjectives, adverbs, and all PoS categories); and (2) 6 expert-aligned narrative similarity scores (BLEU-n, METEOR, ROUGE-L). These reference-based features enable macrostructural analysis by evaluating informativeness and coherence against curated standards. However, as these features derive primarily from word frequencies or local n-gram matching, they are incapable of modeling temporal topic development or the cross-modal consistency.
@@ IV-C DTM-based Temporal Analysis
Figure 1: The topic distribution for (a) the HC group and (b) the NCD group; and the topic evolution for (c) the HC group and (d) the NCD group; as obtained from DTM. Plots (a) and (b) show the weights of the top-10 topic words for each topic (A to E) for the group. Plots (c) and (d) display the average topic proportions with curves and standard deviations with shaded areas. Ordering of topics A to E is based on the positions of their first peaks in topic evolutions.
Temporal changes or topic evolution in narratives are critical for understanding the progression and manifestations of NCDs, as these elements reflect higher-level cognitive abilities. To capture these patterns, we utilize Dynamic Topic Modeling (DTM) [80], which tracks them by explicitly modeling the evolution of topics over time. In contrast to previously mentioned static methods, DTM offers a dynamic perspective on the underlying temporal trends within the narratives.
@@
Dynamic Topic Model (DTM) is an extension of classical topic models. It introduces time dependencies upon the distributions, making it particularly suitable for analyzing temporal changes in narrative content. Formally, DTM models the dynamic nature of topic evolution by introducing the time dependency $t$ in the topic word distribution $\beta_{k,t}$, where $k$ and $t$ denote the topic and time indices respectively. In this way, it allows the topic distributions to drift over time via a Gaussian process, i.e., $\beta_{t}|\beta_{t-1}\sim\mathcal{N}(\beta_{t-1},\boldsymbol{\Sigma})$, where $\boldsymbol{\Sigma}$ represents the covariance matrix controlling topic transition smoothness.
The generative process of DTM proceeds as follows:
First, for each document $d$ in time slice $t$, the model first samples a topic proportion from a Dirichlet prior: $\theta_{d}\sim\text{Dir}(\alpha)$, where $\alpha$ is the Dirichlet concentration parameter controlling the sparsity of topic distributions.
Second, for each word position $n$ in document $d$, the model samples a topic assignment from a multinomial distribution parameterized by the document topic proportion: $z_{d,n}\sim\text{Mult}(\theta_{d})$.
@@
The model parameters are estimated through variational inference [81] or the Expectation-Maximization [82] algorithm during the training phase.
@@
In order to apply the DTM model to narrative analysis for NCD detection, we split each participant’s ASR transcription evenly into 15 slices, assuming that each image leads to approximately the same number of words in spoken narratives. To mitigate fragmented sentences, we shift the split points to the nearest punctuation boundaries. This segmentation allows us to track the evolution of participants’ narrative topics across the story’s progression, enabling a granular analysis of discourse cohesion by observing how thematic content shifts and develops with each sequential image in the narrative.
@@
We apply DTM by inputting the segmented transcription data, where each slice represents a distinct portion of the narrative associated with a time step. The DTM processes this data to identify evolving topics within and across the segmented slices, allowing us to observe how smoothly and coherently topics transition as the story progresses.
@@
Fig. 1 illustrates the topic distribution (a, b) and temporal evolution of topic proportions (c, d) for the HC and NCD groups. For the distribution of topics over words, there exists differences between HC and NCD groups, where HC topics over time tend to be more aligned with the order of the images. For example, the topic words of the HC group, shown in Fig. 1(a), include “home” and “garden” in topic “A” (corresponding to settings of images 1-4), “ant”, “tree hollow” and “goose” in topic “B” (corresponding to images 5-8), “cow” and “stone” in topic “C” (corresponding to images 9-10), “fall” and “pond” in topic “D” (corresponding to images 10-11), and “many” (rabbits) and “mother” in topic “E” (corresponding to images 12-15). These topic words generally align with the different scenes along the story. In contrast, the topic words of the NCD group, shown in Fig. 1(b), are more chaotic. For example, “tree hollow” in topic “B” and “ant, squirrel” in topic “C” both predominantly appear in the same scene in images 5 and 6, indicating overlapping topic assignments.
@@
Another interesting finding is that the HC narratives exhibit a cyclical pattern. Specifically, most HC participants mention “return” and “home” in topic ”E”, even though these elements are not depicted in the final images. This suggests that HCs can tie the ending of the story back to its beginning, completing the girl’s quest to find the lost rabbit and bringing the narrative to a satisfying closure.
@@
For the dynamic proportions of these topics over time, as shown in Fig. 1(c, d), we observe that the HC group exhibits more distinct peaks (with each topic dominating consecutive slices) and smoother transitions between topics over time (peaks rising sharply after the preceding topic declines). In contrast, the NCD group shows more erratic changes with less pronounced peaks (e.g., reduced peak amplitudes and sporadic dominance).
Figure 2: (a) The proposed TITAN model. (b) The Attention module of TITAN. (c) The correlation map between image and text embeddings in TITAN, where the 15 textual chunks are manually segmented and aligned with the corresponding images. TITAN takes images and text narratives as inputs to predict NCDs, with the Attention mechanism using RoPE to emphasize important positions.
@@
Based on these findings, we propose six DTM-related statistics for NCD detection. The first metric, termed “topic consistency”, is derived from the hit rate of the top-10 topic words at each time step, derived from a DTM pretrained on all narratives. This statistic measures how consistently a participant’s topic words align with the top-10 words over time, thus providing insights into the stability and coherence of topics across different groups (HC vs. NCD). For example, if the words in the text slice include all top-10 topic words at the same timestamp $t$, then topic_consistency($t$) = 1.0. Higher topic consistency suggests a more stable thematic structure, indicating a potentially healthier cognitive state. The second metric, termed “topic cycle”, is a Boolean indicator of whether the word “home” appears in the latter half of each narrative.
@@
The remaining four metrics capture the dynamic properties of topic evolution: (1) the standard deviation of topic proportions (topic variability); (2) the temporal correlation of topic proportions over time (topic temporal corr); (3) the peak-to-peak range, defined as the difference between the maximum and minimum values (topic ptp range); and (4) the average rate of change, measured as the Euclidean distance between adjacent topic distributions over time (topic change rate). These statistics capture reflect the development of VSN topics, and can be used as features followed by a downstream classifier for NCD detection, while remaining interpretable through their well-defined mathematical properties.